# A Population Analysis of 20 Exoplanets Observed from the Optical to the Near-infrared Wavelengths with HST: Evidence for Widespread Stellar Contamination

### v2

@@ 5.1 Retrieved Temperature and Water Abundance
In group 1 we identify WASP-121 b Case 1 and 2, WASP-52 b Case 1 and 2, and WASP-6 b Case 2, while group 2 comprises e.g. Case 1 and 2 for HAT-P-11 b, WASP-6 b Case 1, WASP-69 b and WASP-80 b Case 2. These cases highlight that employing a more complex model does not result in the unnecessary adjustment of retrieved parameters. On the other hand, it suggests that the retrieval process remains unbiased towards a more complex parameterisation, particularly in scenarios where stellar activity might not be present. Group 3 is where the majority of the planets lie. Here we place e.g. WASP-69 b, WASP-39 b, HAT-P-12 b Case 1 and GJ-436 b, WASP-29 b, WASP-74 b Case 2. The reason for the discrepancies is found behind model fine-tuning. In Fig. 3 we are just showing how the two simplest models compare, namely the base model with a quiescent star and the base model with an active star, neither of which are completely physically realistic. However, we find that the inclusion of clouds and/or alkali metals on the base model with and without stellar activity considerably improves the fit in most cases and aids in retrieving temperature values closer to the ballpark provided by the calculated equilibrium temperatures. For instance, that is the case for HAT-P-26 b and WASP-127 b. By adding a layer or grey clouds on HAT-P-26 b Case 2, we find a equal to 747 K in the base retrieval and = 1048 K when we consider stellar activity to influence the spectrum (Table E.6). Thus, while both the simplest models result in similar retrieved temperatures, the addition of opaque clouds is crucial to retrieve a more reasonable temperature value with the stellar contamination model w.r.t. to the one without. Likewise, for WASP-127 b (Table E.17), the inclusion of clouds and alkali metal opacities leads to a notable shift in the retrieved temperatures. In the baseline setups, the retrieved temperatures range from 530 to 630 K, whereas in the more complex models—with and without the addition of stellar activity—the temperatures rise to 1100 K and 1520 K respectively. These findings align more closely with the calculated equilibrium temperature of the planet, which is = K.
@@
An analysis of the retrieved water abundances reveals a variety of water volume mixing ratios (VMRs), which are closely related to the spectral slopes and the temperature values. On most planets, Case 1 and 2 spectra often display contrasting optical data, i.e. STIS visits obtained with the same grating at different epochs exhibit distinct slopes with differing gradients. The base model including Rayleigh scattering and water vapour can typically fit a spectrum with a positive slope in the visible wavelengths reasonably well. This is the case of WASP-17 b, WASP-6 b, WASP-69 b and HD 209458 b. The Case 1 spectra of these planets return a between -4 and -6, within the typical range seen for hot-Jupiters. However, an extreme optical slope, that can only be partially described by Rayleigh scattering, requires the inclusion of stellar contamination to model optimally. This is the case of WASP-101 b (Table E.15): its Case 1 base model returns a water VMRs lower than -9 but the addition of stellar contamination brings the value to a nominal range ( -3), hinting that stellar activity may be a relevant source of contamination on this spectrum. Moreover, we find that the base model struggles to fit a spectrum displaying a negative slope in the optical wavelengths. Unocculted stellar faculae, however, provide an plausible explanation for the downward trend: when the planet crosses a transit chord that is characterised by a reduced flux compared to the surrounding stellar surface covered in high-temperature faculae, the resulting signal tends to underestimate the planetary radius. Consequently, this leads to a decrease in transit depth, generating a characteristic negative bluewards slope. In this group, we highlight the case of HAT-P-26 b Case 1 (Table E.6). One of the base model solutions attempts to explain the slope by increasing the water content to a super-solar value ( = ) that is likely unrealistically high. When the stellar contamination model is included a less extreme mixing ratio of is obtained. Similarly for WASP-121 b Case 1 (Table E.16), stellar activity reduces the retrieved water VMR from -1.15 to -6. Considering stellar contamination can therefore eliminate the need for resorting to super-solar metallicity models to explain elevated water values (Welbanks et al., 2019).
@@
It is important to note that the results reported above refer to the baseline models only, i.e. where water is the only active molecular absorber. Water content and planetary temperature are highly correlated. This explains why in a basic scenario with only water absorption, the retrieval can easily find two contrasting solutions: either an excessively high temperature being compensated by a lower water content or a higher water abundance balanced out by a lower planetary temperature. The addition of clouds usually helps to disentangle these degeneracies, largely by reducing the water content and regulating the temperature to a more physically motivated value.
@@ 5.2 Bayes Factor results
The Bayes Factor, as described in Section 4.1 is a crucial criterion to compare how well different model parameterisations describe the data and is by no means specific to stellar activity. Here it provides us with precious insights into how important stellar contamination is to accurately fit a given spectrum. The results of the Bayes Factor calculation, as defined in Equation 5, are reported in Fig. 4 for all planets in the sample, considering, once again, each spectral case separately. The strength of the evidence has been colour-coded according to the Jeffreys’ scale. The scale conventionally applies to the results above zero, but it can be easily mirrored to explain the negative Bayes Factor values. We remind the reader that here a negative value is plausible, as the Bayes Factor, defined as the difference between two logarithms, can yield negative results. For values greater than zero, the stellar contamination model is preferred over the simpler base model. As it is evident from Fig. 4 - Case 1 results, the more complex model is strongly favoured ( 5) for nine out of the eleven spectra for which the stellar activity model is preferred overall. However, HAT-P-26 b and HD 209458 b exhibit no clear preference between the models. If we focus on the lower section of the left-hand side plot, where the Bayes Factor is negative, we observe that among the nine planets for which the Bayesian evidence is in favour of the base model, only GJ 436 b strongly prefers the base model ( = -6.55). Notably this Bayes factor is still significantly smaller compared to those of the majority of planets showing strong preference for the activity model (the largest Bayes factor is seen for WASP-52b where = 26.84). HAT-P-12 b moderately favours the base model, while the remaining spectra (HAT-P-11 b, WASP-127 b, WASP-29 b, WASP-31 b, WASP-39 b, WASP-69 b, WASP-80 b) indicate a weak to inconclusive evidence in favour of the baseline model, suggesting that both the base and the stellar contamination model equivalently explain the data.
@@
Similarly for the Case 2 results, the Bayes Factor of twelve planets lies above zero, ten of which indicate a moderate to strong evidence in favour of stellar activity, while the remaining two indicate an inconclusive evidence or are at the border between a weak and an inconclusive evidence. Although the base model is preferred for six spectra, their evidence is mostly inconclusive or weak, with -2.5 0; even where a moderate evidence is identified for WASP-29 b and WASP-80 b, their values are closer to a weak strength of evidence (-2.5) rather than a strong one (-5). For Case 2, no strong preference for the base model is seen for any of the planets in our sample.
@@
These results clearly indicate that in both Cases the preferences we get for the stellar contamination model are substantially stronger than those we get for the base model. This is in some ways to be expected, as a preference for the base model should predominantly result from the additional dimensionality penalty applied to the contamination model rather than from a goodness-of-fit perspective. In Case 1, only two planets, GJ 436 b and HAT-P-12 b, show a moderate to strong preference for the base model compared to nine planets that strongly or moderately prefer an active star model instead. This distinction is even more evident in the Case 2 results, where six planets strongly prefer the stellar model and none show a preference for the base model stronger that -3. Further evidence supporting the adequacy of the stellar contamination model in describing the spectral data, particularly when a substantial Bayesian evidence favours it, relates to the weighting of the error bars. During the retrieval calculations, the nested sampling algorithm assigns greater weight to smaller error bars, prioritising an optimal fit to the WFC3 data before addressing the STIS data. It is noteworthy that, even though the STIS data are weighted down, the stellar model (which carries the most relevance in the STIS bands) remains the preferred model for nearly half of the spectra in both Case 1 and Case 2 and can also visually improve the fit in the NIR region.

### v3

@@ 3 Spectral Modelling
3.1 The ASteRA retrieval plugin
ASteRA is a stellar activity plugin for TauREx 3.1, designed to account for photospheric heterogeneities in retrievals. This module allows for modelling the host star not only as a homogeneous body characterised by a single temperature and SED but also as an active star composed of multiple temperature components: the quiet photosphere, starspots, and faculae. The spectral emission of the heterogeneous star is modelled using BT-Settl models (Allard et al., 2012), while the stellar spectral grid is generated using the PHOENIX library (Husser et al., 2013). In essence, starspots and faculae are treated as separate cooler and hotter stars, respectively. The stellar emission densities (SEDs) from these three surface components (spots, faculae, and quiet photosphere) are combined based on their covering fractions to produce the observed disk-integrated stellar spectrum:
Here, is the flux of the heterogeneous star at a given wavelength, while , , and represent the spectra of the quiet photosphere, starspots, and faculae, respectively. The parameters and denote the covering fractions of the spots and faculae relative to the observed stellar disk. According to Eq. 4, the heterogeneous star model defines the spot and facula temperatures ( and ), allowing for the calculation of their temperature contrasts relative to the quiescent photosphere. Additionally, the model derives the filling factors for spots and faculae, representing the fraction of the visible stellar disk covered by their projected areas ( and ). As a result, modelling an active star requires introducing two additional fitting parameters (for either spots or faculae) or four (when both are included) compared to a standard retrieval: the spot temperature () and its filling factor (), as well as the facula temperature () and its filling factor ().
ASteRA relies on the fundamental stellar parameters, effective temperature, metallicity, and , to select the corresponding PHOENIX spectra, which are fixed within the retrieval. The positional dependence of spots and their interaction with limb-darkening effects has been thoroughly investigated by Thompson et al. (2024), who found that these effects become significant only in cases of high stellar activity. Given the low-resolution nature of HST observations, we opted for the simplified two-parameter model (temperature and covering fraction) for both spots and faculae.
3.2 Retrievals setup
[…]
where is the host star’s effective temperature, is the stellar radius, is the inverse of the planet’s semi-major axis to stellar radius ratio. When calculating , we set the albedo () and the emissivity () to be 0.2 and 0.8 respectively. The error bars associated with the equilibrium temperatures presented in Table B.1 were computed under the assumption of different albedo and emissivity values. For the lower bound we adopted and , indicating a planet with high reflectivity and radiative properties. For the upper bound, assuming a condition with low reflectivity but a high heat redistribution potentially resulting from the planet’s formation process, we employed and . In all of our retrievals, we modelled the stellar photosphere using the PHOENIX BT-Settl synthetic stellar atmosphere spectra (Allard et al., 2012). We fixed the fundamental stellar parameters, the stellar radii () and masses (), photospheric effective temperatures () and metallicities to literature values, as reported in Table C.1. To account for potential stellar contamination within the retrieval, four additional parameters relating to stellar activity are fit for alongside the planetary parameters of interest. These activity parameters are the spot and facula temperatures, and respectively (assuming that all spots/faculae are characterised by a single temperature), and their corresponding filling factors i.e. coverage as a fraction of the total projected stellar surface, which are denoted as and . The prior bounds on the filling factors and were set between 0.0 and 0.9, where a filling factor of 0.9 would equate to 90% of the photosphere being covered by either of these active regions, with the implicit condition that the sum of these two parameters cannot exceed 1. We set prior bounds for the temperatures of these active regions that are dependent on the effective (photospheric) temperature () of the host star and, as such, their absolute values vary on a case-by-case basis. For the spot temperature, , we set the lower bound to be equal to – 1500K and the upper bound to be equal to . For the facula temperature, , we set the lower bound to be equal to and the upper bound as + 500K. The reasoning behind imposing priors of different magnitudes on the faculae and spot temperatures is that we expect faculae to display a smaller positive temperature contrast () with respect to the photosphere as compared to the larger negative contrast of spots (e.g. Berdyugina, 2005; Gondoin, 2008; Panja et al., 2020; Norris et al., 2023).
4 Stellar Activity metrics
4.1 Bayes factor
To assess the strength of the model preference from a Bayesian perspective, we utilise the Bayes factor, defined as the ratio of two models’ evidences given an observed data set. In our study we compare the evidence of the baseline model to that of the stellar baseline model for each planetary spectra, considering one spectral case at a time. The fitted parameters for the base model include water vapour, , and . Conversely, the stellar model incorporates these three parameters plus four additional activity parameters, as described in Section 3.2. Thus, the dimensionality of the stellar model, i.e. the numbers of parameters in the model, is increased by four. All models used in the Bayes factor calculations are cloud-free and assume a uniform atmospheric temperature and water content with altitude. By employing the simplest atmospheric models we aim to distinguish at first order the signatures that stellar activity imprints on the planetary spectra. The Bayes factor serves as a fully Bayesian approach to model selection, relying solely on the evidence of the respective models. In log space it is expressed as follows:
where represents the Bayesian evidence of the stellar model and corresponds to that of the base model. A negative value suggests a preference for the base model, while a positive value indicates a preference for the model incorporating stellar activity, despite the additional dimensionality it requires. Following the Jeffreys’ scale (Jeffreys, 1998; Trotta, 2008) we establish three thresholds to gauge the extent to which the stellar model is favoured over the baseline model (or vice versa). A Bayes factor 1 2.5 represents a weak evidence against the simpler base model; 2.5 5 is considered to be moderate evidence in favour of including stellar contamination and 5 is considered strong evidence in favour of using a model that incorporates stellar activity. Similarly, the Jeffreys’ scale can be used to explain the preference of the base model over the stellar model when the Bayes factor is negative. For instance, a -5 signifies a strong preference for the base model over the active star model. A Bayes factor falling between -5 and -2.5 suggests a moderate preference for the baseline model, while a ranging from -2.5 to -1 indicates weak evidence for it. The strength of the evidence becomes inconclusive for values 1 -1, effectively implying that both models are equally capable of explaining the observed data. The Bayes factor is a robust indicator for goodness-of-fit and an essential criterion to discriminate against models. Arguably, the greatest strength of the Bayes factor is its tendency to favour simpler models over complex ones in order to prevent over-fitting. This is essential to make sure that our retrieval model is not biased towards finding stellar activity if there is not sufficient evidence for it. Due to the way Bayesian frameworks naturally interact with low-resolution spectra and their associated uncertainties, we explore additional, newly defined metrics. These metrics help provide a broader perspective on the activity environment by examining the contexts in which stellar activity plays a significant role. They are also meant to discern the extent of stellar activity on a qualitative scale, such as distinguishing between low and high activity regimes. A complete overview of these metrics will be given in Thompson et al. (in prep.), where a more extensive description will be presented alongside further applications.
4.2 Stellar Activity Distance metric (SAD)
The Bayesian approach to model selection tends to favour data sets with smaller uncertainties, prioritising precise data during the fitting routine. Consequently, STIS observations, which typically exhibit larger error bars compared to WFC3 due to their lower S/N, will carry less weight during the retrieval process, leading the model to assign less significance to the STIS data points. However, since stellar activity tends to imprint stronger signals within the wavelength range covered by STIS, it is crucial to evaluate how well the retrieval aligns with observations in the visible and to extract as much of the information content of this region as we can. Solely relying on STIS data sets for retrievals is not advisable; meaningful results require anchoring the optical observations to WFC3 data where the bulk of the planetary information content lies. Without the continuum provided by the near-infrared data, the task of constraining basic planetary parameters becomes exceedingly challenging, resulting in larger uncertainties, increased model and parameter degeneracies and a tendency for the retrieval to be pushed towards exploring unphysical territories. Hence, we introduce a metric that, focusing exclusively on the optical regime, aims at evaluating which retrieval model (with or without activity) better aligns with STIS observations without the need for running further retrievals or altering the input observation in any way.
with an uncertainty of
4.3 Stellar Activity Temporal metric (SAT)
To assess the repeatability and consistency of observations for each planet, we examine how the data sets acquired with the same STIS grating compare to each other. We note that stellar activity is not the only phenomenon which could be responsible for spectra taken at different epochs diverging from one another. Other temporally-variable processes e.g. heterogeneous, patchy clouds (MacDonald & Madhusudhan, 2017) or the influence of macro-scale planetary weather/storms (Skinner & Cho, 2022) could also contribute. However, stellar activity is likely the most plausible cause of such strong deviations observed for some of the planets in the sample over these relatively short timescales. This is because the activity modulation is synchronised primarily with the timescales of stellar rotation. This is particularly evident in cases where a reversal in the optical slope is observed between subsequent STIS visits, as seen, for example, in WASP-6 b, WASP-17 b, WASP-80 b, and HD 209458 b. Such reversals may indicate the host star transitioning between regimes dominated by starspots and those dominated by faculae. This metric cannot be used for WASP-127 b and HAT-P-18 b since we possess only one observation per grating for each of these planets. In this context, we disregard retrieval models and concentrate purely on observational data. Our aim is to compare two or more observations of the same planet normalised to a common factor. To maintain uniformity across the entire sample, we use the WFC3 data set of each planet as the normalisation factor for all its STIS observations considered for this metric. For each planet, the normalisation factor is determined as the weighted mean of its WFC3 data set, with weights assigned inversely proportional to the transit depth error bars, thereby giving less weight to larger error bars. Finally, we compute the average of the set of absolute normalised differences per pair of observations ( and ), which we define as the SAT:
For individual planets, determining the average deviation in transit depth in ppm between subsequent STIS visits would be informative on its own. However, to perform comparative planetology we need a method of normalising these differences across our entire population. As each of these planets have atmospheres characterised by different scale heights, for comparability we report the absolute differences as a percentage of the weighted average transit depth of the WFC3 observations, which are available for every planet in our sample. Our assumption here is that at first order, the weighted average transit depth taken from the WFC3 observation is a reasonable proxy for scale height effects, although it will of course be slightly influenced by any absorption features present in that region.
5 Results
5.1 Retrieved temperature and water abundance
[…]
@@ 5.1 Retrieved temperature and water abundance
In group 1 we identify WASP-121 b Case 1 and 2, WASP-52 b Case 1 and 2, and WASP-6 b Case 2, while group 2 comprises e.g. Case 1 and 2 for HAT-P-11 b, WASP-6 b Case 1, WASP-69 b and WASP-80 b Case 2. These scenarios highlight that employing a more complex model does not result in the unnecessary adjustment of retrieved parameters. On the other hand, it suggests that the retrieval process remains unbiased towards a more complex parameterisation, particularly in scenarios where stellar activity might not be present. Group 3 is where the majority of the planets lie. Here we place e.g. WASP-69 b, WASP-39 b, HAT-P-12 b Case 1 and GJ-436 b, WASP-29 b, WASP-74 b Case 2. The reason for the discrepancies is found behind model fine-tuning. In Fig. 3 we are just showing how the two simplest models compare, namely the base model with a quiescent star and the base model with an active star, neither of which are completely physically realistic. However, we find that the inclusion of clouds and/or alkali metals on the base model with and without stellar activity considerably improves the fit in most cases. Furthermore, it aids in retrieving temperature values closer to the ballpark provided by the calculated equilibrium temperatures. For instance, that is the case for HAT-P-26 b and WASP-127 b. By adding a layer or grey clouds on HAT-P-26 b Case 2, we find a equal to 747 K in the base retrieval and = 1048 K when we consider stellar activity to influence the spectrum (Table E.6). Thus, while both the simplest models result in similar retrieved temperatures, the addition of opaque clouds is crucial to retrieve a more reasonable temperature value with the stellar contamination model w.r.t. to the one without. Likewise, for WASP-127 b (Table E.17), the inclusion of clouds and alkali metal opacities leads to a notable shift in the retrieved temperatures. In the baseline setups, the retrieved temperatures range from 530 to 630 K, whereas in the more complex models–with and without the addition of stellar activity–the temperatures rise to 1100 K and 1520 K respectively. These findings align more closely with the calculated equilibrium temperature of the planet, which is = K.
@@
An analysis of the retrieved water abundances reveals a variety of water volume mixing ratios (VMRs), which are closely related to the spectral slopes and the temperature values. On most planets, Case 1 and 2 spectra often display contrasting optical data, i.e. STIS visits obtained with the same grating at different epochs exhibit distinct slopes with differing gradients. The base model including Rayleigh scattering and water vapour can typically fit a spectrum with a positive slope in the visible wavelengths reasonably well. This is the case of WASP-17 b, WASP-6 b, WASP-69 b and HD 209458 b. The Case 1 spectra of these planets return a between -4 and -6, within the typical range seen for hot-Jupiters. However, an extreme optical slope, that can only be partially described by Rayleigh scattering, requires the inclusion of stellar contamination to be modelled optimally. This is the case of WASP-101 b (Table E.15): its Case 1 base model returns a water VMR lower than -9 but the addition of stellar contamination brings the value to a nominal range ( ), hinting that stellar activity may be a relevant source of contamination on this spectrum. Moreover, we find that the base model struggles to fit a spectrum displaying a negative slope in the optical wavelengths. Unocculted stellar faculae, however, provide a plausible explanation for the downward trend: when the planet crosses a transit chord that is characterised by a reduced flux compared to the surrounding stellar surface covered in high-temperature faculae, the resulting signal tends to underestimate the planetary radius. Consequently, this leads to a decrease in transit depth, generating a characteristic negative bluewards slope. In this group, we highlight the case of HAT-P-26 b Case 1 (Table E.6). One of the base model solutions attempts to explain the slope by increasing the water content to a super-solar value ( = ) that is likely unrealistically high. When stellar contamination is included, a less extreme volume mixing ratio of is obtained. Similarly for WASP-121 b Case 1 (Table E.16), stellar activity reduces the retrieved water VMR from to . Considering stellar contamination can therefore eliminate the need for resorting to super-solar metallicity models to explain elevated water values (Welbanks et al., 2019).
@@
It is important to note that the results reported above refer to the baseline models only, i.e. where water is the only active molecular absorber. Water content and planetary temperature are highly correlated. This explains why in a basic scenario with only water absorption, the retrieval can easily find two contrasting solutions: either an excessively high temperature being compensated by a lower water content, or a higher water abundance balanced out by a lower planetary temperature. The addition of clouds usually helps to disentangle these degeneracies, largely by reducing the water content and regulating the temperature to a more physically motivated value.
@@ 5.2 Bayes factor results
The Bayes factor, as described in Section 4.1 is a crucial criterion to compare how well different model parameterisations describe the data and is by no means specific to stellar activity. Here it provides us with precious insights into how important stellar contamination is to accurately fit a given spectrum. The results of the Bayes factor calculation, as defined in Equation 6, are reported in Fig. 4 for all planets in the sample, considering, once again, each spectral case separately. In a few scenarios the retrieval outputs two possible solutions that fit the observations equally well from a statistical perspective. However in Fig. 4 we present either the solution with the highest Bayesian evidence if both models appear physically motivated (e.g. for WASP-17 b Case 1), or the most physically motivated of the two solutions. To define a physically motivated model, we examine the retrieved water abundances and temperature values, comparing them with expectations from the literature. For instance, in the case of GJ 3470 b, the baseline retrievals yield two possible solutions, one of which suggests a very low retrieved temperature (100 K). However, we consider this unphysical due to its inconsistency with existing literature. Therefore, we exclude this solution from further analysis, even though the Bayesian evidence suggests it fits the data better. Another example is HAT-P-26 b. For this planet the baseline retrieval in Case 1 also provides two possible solutions that are generally consistent, except for the water abundance. In the second solution, the retrieved water abundance is nearly two orders of magnitude higher than expected, indicating a super-solar abundance. In this case, we opt for the first solution, as its retrieved water abundance aligns better with other retrievals, and it also has a higher Bayesian evidence.
@@
The strength of the evidence in Fig. 4 has been colour-coded according to the Jeffreys’ scale. The scale conventionally applies to the results above zero, but it can be easily mirrored to explain the negative Bayes factor values. We remind the reader that here a negative value is plausible, as the Bayes factor, defined as the difference between two logarithms, can yield negative results. For values greater than zero, the stellar contamination model is preferred over the simpler base model. As it is evident from Fig. 4 - Case 1 results, the more complex model is strongly favoured ( 5) for nine out of the eleven spectra for which the stellar activity model is preferred overall. However, HAT-P-26 b and HD 209458 b exhibit no clear preference between the models. If we focus on the lower section of the left-hand side plot, where the Bayes factor is negative, we observe that among the nine planets for which the Bayesian evidence is in favour of the base model, only GJ 436 b strongly prefers the base model (). Notably this Bayes factor is still significantly smaller compared to those of the majority of planets showing strong preference for the activity model (the largest Bayes factor is seen for WASP-52b where ). HAT-P-12 b moderately favours the base model, while the remaining spectra (HAT-P-11 b, WASP-127 b, WASP-29 b, WASP-31 b, WASP-39 b, WASP-69 b, WASP-80 b) indicate a weak to inconclusive evidence in favour of the baseline model, suggesting that both the base and the stellar contamination model equivalently explain the data.
@@
Similarly for the Case 2 results, the Bayes factor of twelve planets lies above zero, ten of which indicate a moderate to strong evidence in favour of stellar activity. The remaining two indicate an inconclusive evidence or are at the border between a weak and an inconclusive evidence. Although the base model is preferred for six spectra, their evidence is mostly inconclusive or weak, with -2.5 0; even where a moderate evidence is identified for WASP-29 b and WASP-80 b, their values are closer to a weak strength of evidence (-2.5) rather than a strong one (-5). For Case 2, no strong preference for the base model is seen for any of the planets in our sample.
@@
These results clearly indicate that in both Cases the preferences we obtain for the stellar contamination model are substantially stronger than those we obtain for the base model. This is in some ways to be expected, as a preference for the base model should predominantly result from the additional dimensionality penalty applied to the contamination model rather than from a goodness-of-fit perspective. In Case 1, only two planets, GJ 436 b and HAT-P-12 b, show a moderate to strong preference for the base model compared to nine planets that strongly or moderately prefer an active star model instead. This distinction is even more evident in the Case 2 results, where six planets strongly prefer the stellar model and none show a preference for the base model stronger than .
The weighting of the error bars offers further evidence supporting the adequacy of the stellar contamination model in describing the data, particularly when a substantial Bayesian evidence favours it. During the retrieval calculations, the nested sampling algorithm assigns greater weight to smaller error bars, prioritising an optimal fit to the WFC3 data before addressing the STIS data. It is noteworthy that, even though the STIS data are weighted down, the stellar model (which carries the most relevance in the STIS bands) remains the preferred model for nearly half of the spectra in both Case 1 and Case 2 and can also visually improve the fit in the NIR region.